Grade 12th passGeneral Physics

Time period of an artificial satellite is T of orbit R .What is the radius if it`s time period is 8T

Mahi
9 Years agoGrade 12th pass

1 Answer

ramkumar
9 Years ago
By keplers second  LAW of planetary motion ,T^2=Kr^3
The radius of the satellitete can be given if time period is known,
given, THE SATELLITE has a time period of R then ,T^2=R^3
Then for a satellite which has a time period of 8T the radius of that satellite must be 4R